### v2

# PePNet: A Periodicity-Perceived Workload Prediction Network Supporting Rare Occurrence of Heavy Workload

Abstract
Cloud providers can greatly benefit from accurate workload prediction. However, the workload of cloud servers is highly variable, with occasional heavy workload bursts. This makes workload prediction challenging. There are mainly two categories of workload prediction methods: statistical methods and neural-network-based ones. The former ones rely on strong mathematical assumptions and have reported low accuracy when predicting highly variable workload. The latter ones offer higher overall accuracy, yet they are vulnerable to data imbalance between heavy workload and common one. This impairs the prediction accuracy of neural network-based models on heavy workload. Either the overall inaccuracy of statistic methods or the heavy-workload inaccuracy of neural-network-based models can cause service level agreement violations. Thus, we propose PePNet to improve overall especially heavy workload prediction accuracy. It has two distinctive characteristics: (i) A Periodicity-Perceived Mechanism to detect the existence of periodicity and the length of one period automatically, without any priori knowledge. Furthermore, it fuses periodic information adaptively, which is suitable for periodic, lax periodic and aperiodic time series. (ii) An Achilles’ Heel Loss Function iteratively optimizing the most under-fitting part in predicting sequence for each step, which evidently improves the prediction accuracy of heavy load. Extensive experiments conducted on Alibaba2018, SMD dataset and Dinda’s dataset demonstrate that PePNet improves MSE for overall workload by 11.8% on average, compared with state-of-the-art methods. Especially, PePNet improves MSE for heavy workload by 21.0% on average.
Index Terms:
Time series forecasting, cloud computation.
I Introduction
Accurate workload prediction brings huge economic benefits to cloud providers [1] and many cloud frameworks make real-time adjustments based on the workload prediction results [2]. On the one hand, workload prediction provides meaningful insights to improve the utilization of cloud servers while ensuring the quality of service (QoS) [3] [4]. On the other hand, it can also alarm the forthcoming uncommon heavy workload, thus helping to avoid service level agreement (SLA) violations.
(a) The overall and heavy-workload prediction accuracy.
(b) Prediction of GRU on Alibaba2018.
(c) The first half of Achilles’ Heel Loss Function fitting process.
(d) The second half of Achilles’ Heel Loss Function fitting process.
Figure 1: (a)-(b) Overall workload and heavy workload prediction accuracy. (c)-(d) The fitting process of Achilles’ Heel Loss Function. (c) In the first half of the process, the most under-fitting part is on the right. (d) After some training iterations, the prediction error on the right diminishes and the most under-fitting part shifts to the left. Then, Achilles’ Heel Loss Function pays more attention to the left part in this phase.
However, the high variability of cloud-server workload [5] [6] makes workload prediction challenging. Statistical methods and neural-network-based methods are two mainstream workload prediction methods. The former ones require the time series to satisfy strong mathematical assumptions and are not robust when predicting highly variable workload [5]. For example, the classical statistic method, ARIMA [7], requires the time series to be stationary after difference and shows dissatisfied results when predicting highly variable workload [5]. The neural-network-based methods are more suitable to predict highly variable workloads, but they are vulnerable to data imbalance between heavy workload and common one (i.e. the heavy workload is much more uncommon [8]). To demonstrate this, we present statistics of the heavy workload proportion of different datasets in Tab. I, where the heavy workload is defined as the workload greater than the average workload plus one standard deviation for each machine. It has also been proven that the data imbalance between heavy workload and common workload impairs the former’s accuracy [9]. As an empirical example, Fig. 1(a)-Fig. 1(b) shows the overall prediction accuracy and heavy-workload prediction accuracy of some prevailing methods (ARIMA [7], LSTM with attention (LSTMa) [10], Informer [11], Learning based Prediction Algorithm for cloud workload (L-PAW) [5], LSTM and GRU) on Alibaba2018 dataset, where the heavy-workload predicted error is nearly twice as large as the overall predicted error. The inaccurate prediction will not only reduce the utilization of cloud servers but also bring SLA violations, as it provides wrong information to the scheduler.
TABLE I: Proportion of heavy workload.
Dataset Heavy workload proportion
CPU usage of Alibaba’s dataset [^1] https://github.com/alibaba/clusterdata 15.73%
Memory usage of Alibaba’s dataset 12.89%
Workload of Dinda’s dataset [12] 11.15%
To predict overall workload and heavy workload accurately, we propose a Periodic-Perceived Workload Prediction Network (PePNet), which supports the prediction of highly variable workload with the rarely-occurring heavy workload. PePNet mainly improves the prediction accuracy in two aspects: 1) using the periodically recurrent patterns to guide the workload prediction and 2) using an Achilles’ Heel Loss Function [^2] Achilles’ Heel originates from Greek mythology and is the fatal weakness of the hero Achilles, which pays more attention to the under-fitting part to offset the negative influence of data imbalance.
The existing methods that embed the periodic information [13, 14, 15, 16] are applied to the situations where the periodicity is stable and known (i.e. whether the time series is periodic and the length of a period are fixed and known). However, the cloud environment is complex and highly dynamic. Different servers exhibit different periodicity and even the same server exhibits different periodicity at different time. This calls for a technique that detects periodicity and period length automatically and is capable of manipulating time series with mixed periodicity (i.e. the time series is periodic for some time and aperiodic for the other time). Thus, we propose a Periodicity-Perceived Mechanism, which consists of Periodicity-Mining Module and Periodicity-Fusing Module. The Periodicity-Mining Module detects the periodicity and period length and the Periodicity-Fusing Module adaptively fuses the periodic information for periodic and aperiodic time series. We mathematically prove the error bound of the periodic information extracted by Periodicity-Mining Module. Furthermore, we design an automatic hyperparameter determination method to provide convenience for the training process.
The Achilles’ Heel Loss Function specifically targets the upper bound of the prediction error, which enables PePNet to strengthen and improve the most vulnerable part of the prediction. At each training step, the Achilles’ Heel Loss Function picks the most under-fitting part in a predicting sequence and minimizes its prediction error. As shown in Fig.1(c)-1(d), when a kind of pattern is well learned, the most under-fitting part will shift to the position of another pattern. Thus, each part in the predicting sequence is fitted alternatively. This characteristic helps alleviate the problem of overemphasizing sporadic patterns, a limitation that has been observed in many loss functions designed to combat under-fitting issues [17]. In this way, Achilles’ Heel Loss Function solves the under-fitting problem of heavy load and improves the heavy-load prediction accuracy as well as the overall prediction accuracy.
The main contributions of our work are summarized as:
  • We design a Periodicity-Perceived Mechanism, which mines and adaptively fuses the periodic information. Moreover, we mathematically prove an error bound of periodic information extracted by Periodicity-Mining Module and propose an automatic hyperparameter determining method for it.
  • We have formulated an Achilles’ Heel loss function aimed at enhancing the accuracy of heavy workload prediction and overall prediction performance. This approach effectively mitigates the issue of overemphasizing sporadic patterns, a common limitation faced by many loss functions designed to combat under-fitting problems.
  • We conduct extensive experiments on the Alibaba2018, SMD and Dinda’s datasets and demonstrate that PePNet improves the heavy-workload prediction accuracy (MSE) by 21.0% on average, while promoting overall prediction accuracy (MSE) by 11.8% on average.
Figure 2: The overview of PePNet.
II Methodology
Preliminary. The input of data $X$ is divided into three parts, which are denoted by $X^{enc}_{short}\in\mathbb{R}^{I\times d}$, $X^{enc}_{long}\in\mathbb{R}^{M\times d}$, $X_{period}\in\mathbb{R}^{P\times d}$ respectively. The $d$ is the dimension of the feature at each time slot and $I,M$ and $P$ respectively stand for the length of short-term dependent information, long-term tendency information and periodic information. The data division is shown in Fig. 3(a): $Y$ is the ground truth value of predicting workload; $X^{enc}_{short}$ is the nearest workload time series before the predicting part and shows high relevance to predicting workload; $X^{enc}_{long}$ is a bit longer workload time series before $X^{enc}_{short}$ and reflects the long term tendency in workload variation; $X_{period}$ is the first period of workload for each machine; The $y_{period}$ is the workload sequence corresponding to $Y$ in the period of $X_{period}$. The extracting process of $X_{period}$ and $y_{period}$ is illustrated in section II-B1. We use $X^{enc}_{long,i}$ and $X^{enc}_{short,i}$ to denote the $i$-th time slot’s workload in $X^{enc}_{long}$ and $X^{enc}_{short}$ respectively.
The three kinds of information (i.e. short-term dependent information, long-term tendency information and periodic information) bring different effects on workload prediction. The short-term dependent information $X^{enc}_{short}$ is highly related to the predicting sequence, as the customers’ behaviors are continuous and the workload is similar in a short time slot. The long-term tendency information denotes the forthcoming heavy load, as the heavy load is more likely to happen in ascending tendency. This tendency information is usually masked by noises in the short term and has to be extracted from long-term time series. As shown in Fig. 3(b), we pick a small segment of the workload from a long-term-ascending sequence. Though the long-term ascending tendency it is, the short-term workload just fluctuates around a stable value and does not show any increasing tendency. The periodic information provides the recurrent patterns in history and promotes both overall prediction accuracy and heavy-load prediction accuracy.
(a) The input data division.
(b) The long-term tendency information
(c) Local periodicity violation in lax periodic series.
(d) Autocorrelation coefficient for periodic sequence.
(e) Autocorrelation coefficient for aperiodic sequence.
(f) The method of computing periodic information.
(g) Histogram of values for the first peak of the autocorrelation coefficient.
(h) How the value of $\gamma$ impacts the distribution of gradient weights.
Figure 3: Illustrations for data division and periodicity-perceived mechanism.
II-A Overview of PePNet
PePNet is based on an encoder-decoder [18] architecture and fuses three kinds of information from rough data: the long-term tendency information, the short-term dependent information and the periodic information. Besides, the whole model is trained by an Achilles’ Heel Loss Function to offset the negative impact of data imbalance. The overview of PePNet is shown in Fig. 2.
In the encoder, PePNet uses a backbone encoder, which can be implemented as many popular time-series-manipulating structures, such as transformer [19], LSTM [20] etc., to capture short-term dependent information from $X^{enc}_{short}$ and uses Periodicity-Mining Module to detect and extract periodic information from $X_{period}$. As for long-term tendency information, PePNet firstly downsamples $X^{enc}_{long}$, as it helps to improve the processing efficiency as well as maintain and magnify the trend information of time series. Then, PePNet uses self-attention to manipulate the down-sampled sequence. The processes of extracting short-term, long-term and periodic information are shown in Eq.1-3. In function $\operatorname{Attention}(q,k,v)$, $q$, $k$, $v$ respectively stand for the query, key and value[21].
$\overline{X}^{enc}_{short,i+1}=\operatorname{Backbone-Encoder}(X^{enc}_{short,i})$ (1)
$y_{period},X_{period}=\operatorname{PeriodicityMining}(X,X^{enc}_{short})$ (2)
$\overline{X}^{enc}_{long}=\operatorname{Downsample}(\operatorname{Attention}(X^{enc}_{long},X^{enc}_{long},X^{enc}_{long}))$ (3)
In the decoder, PePNet uses a backbone-decoder, which can be implemented as many popular time-series-manipulating structures, such as transformer, LSTM, etc., to generate $\hat{y}\in\mathbb{R}^{J\times d}$, where $J$ stands for the predicting length. When using recurrent neural networks such as LSTM or GRU, the concatenation of $h_{j}$ in Eq. 4 serves as the hidden state, and $\hat{y}_{j}$ serves as the network input. When using a transformer or its variants, $\hat{y}_{j}$ serves as the query and $h_{j}$ serves as both the keys and values. After that, PePNet uses Periodicity-Fusing Module to estimate the reliability of $\hat{y}$ and fuses it with $y_{period}$ according to their reliability. By then, the Periodicity-Fusing Module generates the final prediction $y$. The process of decoder is shown in Eq.4-6, where $W_{h},W_{c}$ are both model parameters.
$\hat{y}_{j+1}=\operatorname{Backbone-Decoder}(\hat{y}_{j},h)$ (4)
$h=\operatorname{Attention}(X^{enc}_{short},\overline{X}^{enc}_{long},\overline{X}^{enc}_{long})$ (5)
$y=\operatorname{PeriodicityFusing}(X^{enc}_{short},y_{period},\hat{y})$ (6)
When training the model, the imbalance between heavy load and normal workload usually leads to unsatisfactory accuracy of heavy load prediction. Thus, we propose an Achilles’ Heel Loss function, which aims to minimize the upper bound of prediction error. At each step, Achilles’ Heel Loss Function iteratively optimizes the most under-fitting part for this step, in order to improve the prediction accuracy of the heavy load when keeping the overall prediction accurate. It is named after Achilles’ Heel as it focuses on the most vulnerable portion along the predicting sequence. Thus, the prediction accuracy of the heavy load becomes more accurate, when keeping the overall prediction accurate.
II-B Periodicity-Perceived Mechanism
Periodic information can effectively improve the overall prediction accuracy as well as the heavy workload prediction accuracy. It consists of Periodicity-Mining Module and Periodicity-Fusing Module.
There are three main challenges to fuse periodic information. (1) We have no priori knowledge about the periodicity of workload for different machines, which calls for the Periodicity-Mining Module to detect the existence of periodicity and the length of one period. (2) The periodicity for different machines is variable (i.e., strict periodic, lax periodic, aperiodic), which calls for an adaptive Periodicity-Fusing Module. (3) It is hard to fuse the periodic information in lax periodic series because lax periodic series are sometimes periodic and sometimes not. In Fig. 3(c), we overlap a series shifted one period ahead and the original series. As Fig. 3(c) shows, there are mainly two obstacles to fuse periodic information in lax periodic series: periodic shift and local periodicity violation. The former one can be solved by dynamic matching which is depicted in section Periodicity-Mining Module. The latter is caused by noise and external events. Therefore, we filter out the noise in periodic information in Periodicity-Mining Module and evaluate the reliability of periodic information in the Periodicity-Fusing Module.
II-B1 Periodicity-Mining Module
PePNet calculates the time series autocorrelation coefficient $\rho_{k}$ for each machine as shown in Eq.7, which represents the linear correlation among all workloads with interval $k$ (i.e., linear correlation between $X_{t}$ and $X_{t-k}$, for every $t\in\{t\in\mathbb{N}^{+}|t<sequence\ length\}$, $X_{t}$ stands for the workload at $t$-th time slot). As shown in Fig. 3(d), when the workload is periodic, the autocorrelation coefficient rises to a large value again after the first decline, which represents a high hop relevance of workload. While the sequence is aperiodic, the autocorrelation coefficients are shown in Fig. 3(e), which have a distinct pattern. Therefore, PePNet sets a hyperparameter $\mathcal{T}$ and judges whether a time series is periodic by detecting whether the autocorrelation coefficient crosses over $\mathcal{T}$ again. The value of $\mathcal{T}$ denotes the acceptable range of periodicity strictness.
$\rho_{k}=\frac{cov(X_{t-k},X_{t})}{\sqrt{cov(X_{t},X_{t})cov(X_{t-k},X_{t-k})}}$ (7)
The position of the first peak in autocorrelation coefficients of periodic sequence strikes the highest hop relevance and denotes the length of one period. We illustrate it in Fig. 3(d). The reason why the autocorrelation coefficient shows such a pattern is that the user behavior is continuous in time, so the autocorrelation coefficients of small $k$ are large. As the time interval $k$ increases, the relevance drops down. But when the time interval $k$ reaches around the integer multiple of period length, the workload at the two moments $X_{t}$ and $X_{t-k}$ shows a high linear correlation again.
Based on these observations, as shown in Algorithm 1, PePNet traverses the value of $k$ from small to large. If PePNet finds the smallest $k$ that makes the autocorrelation coefficient $\rho_{k}$ satisfy the conditions: $\rho_{k}>\rho_{k-1},\rho_{k}>\rho_{k+1}$ and $\rho_{k}>\mathcal{T}$, the time series is periodic and the length of one period is $k$. Otherwise, the time series is aperiodic. We cut out the first period of the machines whose workloads are periodic in the training set as a periodic information knowledge base, which is denoted by $X_{period}$. To solve the issue of period shift, PePNet uses a dynamic matching. If the workload is periodic, PePNet finds the sequence $\{X_{t_{a}},\ldots,X_{t_{a}+I-1}\}$ in $X_{period}$, whose distance from $X^{enc}_{short}$ is the smallest. The distance can be computed either by Mean Square Error (MSE) or by Dynamic Time Warping (DTW). Then, PePNet sets $y_{period}=\{X_{t_{a}+I},\ldots,X_{t_{a}+I+J}\}$. Otherwise, $\{-1,-1,\ldots,-1\}$ is set to $y_{period}$. Such a process is illustrated in Fig. 3(f).
Data: Workload time series $X$, Short-term dependent information $X^{enc}_{short}$
Result: The periodic information for prediction $y_{period}$, The periodic knowledge base $X_{period}$
1 $k\leftarrow-1$;
2 while $k<length(X)$ do
3      $k\leftarrow k+1$;
4       Compute $\rho_{k}$;
5       if $\rho_{k}>\rho_{k-1}\ and\ \rho_{k}>\rho_{k+1}\ and\ \rho_{k}>\mathcal{T}$ then
6            Period length=k ;
7             FoundPeriodLength=True ;
8             break ;
9       end if
10
11 end while
12if FoundPeriodLength then
13       $X_{period}=X[:Period\ length]$;
14       $y_{pos}=\arg\min_{t_{a}}Distance(X[t_{a}:t_{a}+I],X^{enc}_{short})$;
15       $y_{period}=X[y_{pos}:y_{pos}+I]$;
16
17 end if
18else
19       $X_{period}=-1$;
20       $y_{period}=[-1,\ldots,-1]$;
21
22 end if
Algorithm 1 Periodicity-Mining Module
Error estimation. We use $E[(X_{t-k}-X_{t})^{2}]$ to measure the quality of periodic length $k$ found by the Periodicity-Mining Module. We have the following result, where $\sigma_{t-k},\sigma_{t},\mu_{t-k},\mu_{t}$ are the standard deviation and expectation at $t-k$ and $t$ time slots respectively. The $\Delta\sigma$ and $\Delta\mu$ are $\sigma_{t-k}-\sigma_{t}$ and $\mu_{t-k}-\mu_{t}$ respectively. $\sigma$ is the standard deviation of a stationary time series.
Theorem 1. When using Algorithm 1, $E[(X_{t-k}-X_{t})^{2}]$ is upper bounded by $(\Delta\sigma)^{2}+(\Delta\mu)^{2}+2(1-\mathcal{T})\sigma_{t-k}\sigma_{t}$. Furthermore, when the time series is stationary, $E[(X_{t-k}-X_{t})^{2}]$ is upper bounded by $2(1-\mathcal{T})\sigma^{2}$.
Proof of Theorem 1. According to algorithm 1, when PePNet finds $k$, Eq.8-9 holds, where $a_{i}=X_{i}-\mu_{i}$, $\tilde{t}=t-k$. $E[(X_{\tilde{t}}-X_{t})^{2}]$ can be transformed to Eq.10. Then, we substitute Eq.9 into Eq.10 and obtain Eq.11. The Eq.11 can be further reduced into Eq.12 and the first part of Theorem 1 is proven. When the time series is stationary, $\Delta\sigma^{2}$ and $\Delta\mu^{2}$ are zero. Then the upper bound can be reduced to $2(1-\mathcal{T})\sigma_{t-k}\sigma_{t}$.
$\frac{E[a_{\tilde{t}}a_{t}]}{\sigma_{\tilde{t}}\sigma_{t}}>\mathcal{T}$ (8)
$E[a_{\tilde{t}}a_{t}]>\mathcal{T}*\sigma_{\tilde{t}}\sigma_{t}$ (9)
$E[(X_{\tilde{t}}-X_{t})^{2}]=E[a_{\tilde{t}}^{2}+a_{t}^{2}-2a_{\tilde{t}}a_{t}+2\Delta\mu\Delta a+\Delta\mu^{2}]$ (10)
$E[(X_{\tilde{t}}-X_{t})^{2}]<E[a_{\tilde{t}}^{2}]+E[a_{t}^{2}]-2*\mathcal{T}\sqrt{E[a_{\tilde{t}}^{2}]E[a_{t}^{2}]}+\Delta\mu^{2}$ (11)
$E[(X_{\tilde{t}}-X_{t})^{2}]<\Delta\sigma^{2}+\Delta\mu^{2}+2(1-\mathcal{T})\sigma_{\tilde{t}}\sigma_{t}$ (12)
Taking a further look at the upper bound, the first and second items of the upper bound are constants. Thus, the value of error depends on the last item, which is determined by the value of $\mathcal{T}$. Considering that the $\mathcal{T}$ is less than or equal to one, the closer the $\mathcal{T}$ is to one, the smaller the error is. However, most of the time series are lax periodic in reality. Setting the $\mathcal{T}$ as 1 may miss many lax periodic time series. Thus, there is a tradeoff between improving the accuracy of periodic length and finding all the periodic time series (including lax periodic time series).
Considering the tradeoff may make it challenging to choose a proper value of $\mathcal{T}$, we provide an automatic setting method for $\mathcal{T}$ based on statistical observations. For finer tuning, this method provides a searching range for $\mathcal{P}$. The hyperparameter $\mathcal{P}$ is a threshold setting for the first peak of the autocorrelation coefficient. Thus, we look into the distributions of the first peak value. Taking Alibaba2018-CPU [^3] https://github.com/alibaba/clusterdata as an example, we take 75 machines in it and plot a histogram of first-peak value in Fig. 3(g). There are two properties of the distribution: clustered and following Gaussian Distribution for each cluster. We use the Gaussian Mixture Model (GMM) algorithm to fit the mixed Gaussian distribution of its first peak values and plot the Gaussian distributed curve on the histogram in Fig. 3(g). The parameter for this GMM is given in Tab. II, where ”std” is the shorthand for standard deviation. Different clusters have different degrees of periodicity. The higher the expectation of the cluster is, the stricter its periodicity is. To guarantee the quality of the periodicity the Periodicity-Mining Module detects, it picks the workload in the cluster with the highest expected value as periodic and filters the others. Let $\mu$ and $\sigma$ denote the expectation and standard deviation of the cluster with the highest expectation. In industrial applications, $\mu-\sigma$ can be directly used as the $\mathcal{P}$ to save the labor costs, as the probability that the first peak is greater than $\mu-\sigma$ in this cluster is 84.2% according to Gaussian distribution. For finer tuning, the $mathcal{P}$ can be explored between $\mu-\sigma$ and $\mu+\sigma$.
TABLE II: Parameters of Gaussian distribution.
weight expectation std
Gaussian distribution 1 0.84 0.53 0.06
Gaussian distribution 2 0.10 0.23 0.03
Gaussian distribution 3 0.06 -0.04 0.05
II-B2 Periodicity-Fusing Module
To filter out the random noise in the periodic information, PePNet firstly uses an auto-encoder [22]. To solve the problem of variable periodicity for different machines and the problem of local periodicity violation, PePNet uses an attention mechanism to evaluate the reliability of periodic information. The final prediction $y$ is given in Eq.13-14.
$y_{period}^{{}^{\prime}}=Autoencoder(y_{period})$ (13)
$y=Attention(X^{enc}_{short},(\hat{y},y_{period}^{{}^{\prime}}),(\hat{y},y_{period}^{{}^{\prime}}))$ (14)
II-C Achilles’ Heel Loss Function
Heavy load rarely occurs, which leads to poor performance of heavy load prediction due to data imbalance. To solve this problem, PePNet deploys the Achilles’ Heel Loss Function. As analyzed below, the Achilles’ Heel Loss Function actually forces the loss to descend along the gradient directions of some of the largest prediction errors. When iteratively training the model, the Achilles’ Heel Loss Function optimizes the most under-fitting part in one step. In the next step, the original largest prediction error diminishes and the most under-fitting part shifts to another place. In this way, the Achilles’ Heel Loss Function optimizes each part of the sequence iteratively and effectively improves the accuracy of heavy load, when keeping high predicting accuracy of overall workload. The Achilles’ Heel Loss Function is given in Eq.15, where $T$ equals $\{t_{1},\ldots,t_{1+J}\}$, which are the time slots of a forecasting sequence. $y_{t}$ denotes the predicted workload at time $t$ (the prediction length for one sample should be at least 2), $Y_{t}$ denotes the ground-truth workload at time $t$, and $\gamma$ is a hyperparameter.
$l(Y,y)=\gamma\log(\sum_{t\in T}exp(\frac{(y_{t}-Y_{t})^{2}}{\gamma})),\gamma>0$ (15)
The Achilles’ Heel Loss Function actually distributes more weight to the gradient for higher prediction error when back propagating. This aspect is obvious by deriving Eq.15, as shown in Eq.16, where $\alpha_{t}=\frac{exp{\frac{(y_{t}(\mathcal{P})-Y_{t})^{2}}{\gamma}}}{\sum_{t=1}^{T}\exp{\frac{(y_{t}(\mathcal{P})-Y_{t})^{2}}{\gamma}}}$, $\mathcal{P}$ stands for the parameters of PePNet. $\alpha_{t}$ is actually a normalized weight, whose value depends on the prediction error at time slot $t$. $\gamma$ is a scale factor. When $\gamma$ is large, the difference in prediction errors across time slots is narrowed and the weights are relatively uniform for each time slot. When the $\gamma$ is small, the difference in prediction errors across time slots is amplified and the weights become polarized. To visually demonstrate this point, we plot the value of $\alpha_{t}$ for different combinations of $\gamma$ and prediction error $y_{t}-Y_{t}$ in Fig. 3(h). In Fig. 3(h), each row shows the value of $\alpha_{t}$ for a set of prediction error $\{y_{t}-Y_{t}|0\leq t\leq 9\}$, when specifying the $\gamma$. As Fig. 3(h) shown, for a specific $\gamma$, the larger the prediction error is, the bigger the $\alpha_{t}$ is. Besides, for a specific prediction error in a specific error set, the smaller the $\gamma$ is, the bigger the $\alpha_{t}$ is. When $\gamma$ is set to 0.1, $\alpha_{t}$ for the largest prediction error in this row approximates 1.
$\frac{\partial l(Y,y)}{\partial\mathcal{P}}=\sum_{t=1}^{T}(\alpha_{t}\cdot\frac{\partial{(y_{t}(\mathcal{P})-Y_{t})^{2}}}{\partial\mathcal{P}})$ (16)
It is worth noticing that when the $\gamma$ is indefinitely close to 0, the Achilles’ Heel Loss Function becomes a simple formation, which is given in the following theorem. The theorem implies that when $\gamma$ indefinitely approximates 0, only the gradient for maximum prediction error is assigned a weight of 1, while others are assigned a weight of 0.
Theorem 2. As the $\gamma$ is infinitely close to 0, the gradient of the loss function infinitely approximates Eq.17. The $t^{*}=\mathop{\arg\max}_{t}(y_{t}-Y_{t})^{2}$ and the $g(t^{*},\mathcal{P})=(y[t^{*}]-Y[t^{*}])^{2}$.
$\frac{\partial l(Y,y)}{\partial\mathcal{P}}=\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}$ (17)
Proof of Theorem 2. The derivation of $l(Y,y)$ is shown in Eq.18. Then, both the numerator and dominator on the right-hand side are divided by $\exp(\frac{g(t^{*},\mathcal{P})}{\gamma})$ simultaneously and the first step of Eq.19 is obtained. The $g(t^{*},\mathcal{P})$ is surely bigger than $g(t,\mathcal{P}),\forall t\neq t^{*}$, as the definition of $t^{*}$ suggests. Thus, $g(t,\mathcal{P})-g(t^{*},\mathcal{P})<0,\forall t\neq t^{*}$. When the $\gamma(\gamma>0)$ infinitely approaches zero, $exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})$ will also infinitely approaches zero. In this way, the second step of Eq.19 is obtained.
$\frac{\partial l(Y,y)}{\partial\mathcal{P}}=\frac{\sum_{t}^{T}[exp(\frac{g(t,\mathcal{P})}{\gamma})\frac{\partial g(t,\mathcal{P})}{\partial\mathcal{P}}]}{\sum_{t}^{T}exp(\frac{g(t,\mathcal{P})}{\gamma})}$ (18)
$\mathop{\lim}_{\gamma\rightarrow 0}\frac{\partial l(Y,y)}{\partial\mathcal{P}}$ (19)
$=\mathop{\lim}_{\gamma\rightarrow 0}\frac{\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}+\sum_{t,t\neq t^{*}}^{T}[exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})\frac{\partial g(t,\mathcal{P})}{\partial\mathcal{P}}]}{1+\sum_{t,t\neq t^{*}}^{T}exp(\frac{g(t,\mathcal{P})-g(t^{*},\mathcal{P})}{\gamma})}$
$=\frac{\partial g(t^{*},\mathcal{P})}{\partial\mathcal{P}}$
III Experiment
In this section, we conduct extensive experiments to validate the following findings:
  • PePNet improves the accuracy of heavy-workload prediction as well as the overall prediction accuracy compared with the state-of-the-art methods.
  • PePNet only introduces slight time overhead for training and inference.
  • PePNet is insensitive to hyperparameters and has high robustness robustness.
  • Each mechanism in PePNet is proven to play an important role by using ablation experiments.
  • Due to its adaptability for periodicity, when the time series is aperiodic, PePNet still works well.
TABLE III: Basic hyperparameters of PePNet.
Hyperparameter Value
Batchsize of Alibaba2018-CPU 100
Learning rate of Alibaba2018 0.001
Learnig rate of Dinda’s dataset 0.0005
Learning rate of SMD 0.01
$\mathcal{T}$ of Alibaba2018-Mem 0.2
$\mathcal{T}$ of others 0.5
$\gamma$ $\mathop{\lim}_{\gamma\rightarrow 0}$
Number of layers of LSTM 2
Input length of Alibaba2018 for short-term information 50
Input length of Alibaba2018 for long-term information 100
Input length of Dinda’s dataset for short-term information 30
Input length of Dinda’s dataset for long-term information 80
Input length of SMD for short-term information 50
Input length of SMD for long-term information 100
Hidden layer size of LSTM 80
Predicted length 2
Init downsampling step size 20
Number of layers of auto-encoder 1
Number of layers of auto-decoder 1
Output size of auto-encoder 1
Sliding window length for downsampling 100
III-A Experiment Setup
Hyperparameters. We summarize the most important hyperparameters of PePNet in Tab. III.
Baseline Methods. We compare PePNet with the state-of-the-art time series predicting methods and popular workload predicting models and the variants of PePNet.
  • Autoformer (Autof) [23]: Autoformer is a novel decomposition architecture with an Auto-Correlation mechanism. It breaks the pre-processing convention of series decomposition and renovates it as a basic inner block of deep models. Further, Autoformer also has an Auto-Correlation mechanism based on the series periodicity.
  • LSTM with attention mechanism (LSTMa) [10]: The approach extracts the sequential and contextual features of the historical workload data by an encoder network and integrates an attention mechanism into the decoder network.
  • Informer (Inf) [11]: The method uses a ProbSparse attention mechanism and a halving cascading layer to extract the input information and use a generative style decoder to predict. Informer is one of the most recognized methods of time series prediction in recent years.
  • L-PAW [5]: The method integrates top-sparse auto-encoder (TSA) and gated recurrent unit (GRU) block into RNN to achieve the adaptive and accurate prediction for highly-variable workloads. L-PAW is a model specifically designed for workload prediction and is also widely used.
  • LSTM [20]: A classic time series processing model, which uses the hidden state to store long sequences of information.
  • GRU [18]: A variant of the LSTM that modifies the forget gate structure in LSTM to make the model simpler.
  • Reformer (Ref) [24]: Reformer introduces two techniques to improve the efficiency of Transformers. Firstly, it replaces dot-product attention with one that uses locality-sensitive hashing. Secondly, it uses reversible residual layers instead of standard residuals to reduce the memory consumption of the model. It is a widely recognized method in time series prediction.
  • Fedformer (Fedf) [25]: Fedformer combines Transformer with a seasonal-trend decomposition method, in which the decomposition method captures the global profile of time series while Transformers capture more detailed structures. Fedformer is an effective and novel method in time series prediction.
  • Variants of PePNet: We implement the backbone-encoder and backbone-decoder as transformer and LSTM respectively and denote them as PePNet-I and PePNet-II. Without specification, we use PePNet to denote PePNet-II in the following. Moreover, we also introduce some variants for ablation study. We use PePNet- to denote PePNet-II that removes the Periodicity-Perceived Mechanism. We use PePNet to denote the PePNet-II that uses MSE as a loss function. We use PePNet to denote the PePNet-II that both removes the Periodicity-Perceived Mechanism and does not use the Achilles’ Heel loss function.
TABLE IV: The overall and heavy-workload prediction accuracy on three datasets. The MSE in the table are multiplied by 1000.
LSTM LSTMa GRU Inf L-PAW Ref Autof Fedf PePNet- PePNet PePNet PePNet-I PePNet-II
MAPE 0.093 0.117 0.108 0.100 0.102 0.104 0.211 0.110 0.089 0.087 0.116 0.047 0.076
Overall MSE 5.515 8.233 9.516 4.865 10.960 8.950 37.636 9.211 5.773 6.146 6.847 4.420 3.697
Dinda’s MAE 0.048 0.071 0.070 0.041 0.078 0.053 0.149 0.057 0.046 0.049 0.059 0.054 0.029
Dataset MAPE 0.051 0.079 0.094 0.030 0.046 0.040 0.079 0.032 0.035 0.043 0.058 0.065 0.018
Heavy-Workload MSE 12.260 17.925 24.191 6.837 15.853 9.202 23.817 8.729 5.635 7.325 13.274 8.760 6.124
MAE 0.068 0.104 0.123 0.040 0.058 0.051 0.097 0.041 0.046 0.057 0.076 0.083 0.025
MAPE 0.179 0.238 0.287 0.237 0.296 0.278 0.488 0.290 0.207 0.191 0.161 0.076 0.154
Overall MSE 3.434 3.627 3.889 3.282 3.909 4.950 8.021 2.350 3.144 2.974 3.101 0.552 1.757
SMD MAE 0.019 0.023 0.025 0.022 0.025 0.028 0.046 0.023 0.020 0.019 0.018 0.015 0.014
Dataset MAPE 0.150 0.163 0.154 0.166 0.154 0.316 0.367 0.157 0.138 0.129 0.158 0.204 0.112
Heavy-Workload MSE 10.437 10.622 11.945 11.024 11.912 17.587 27.465 5.531 10.941 9.751 10.399 4.138 6.729
MAE 0.037 0.045 0.046 0.041 0.046 0.068 0.095 0.034 0.042 0.038 0.040 0.055 0.033
MAPE 0.039 0.095 0.049 0.058 0.157 0.086 0.037 0.025 0.050 0.027 0.126 0.017 0.016
Overall MSE 0.375 0.489 0.462 0.408 12.721 0.880 1.173 0.510 0.359 0.341 0.975 0.410 0.323
Alibaba MAE 0.014 0.017 0.015 0.015 0.058 0.023 0.026 0.016 0.014 0.013 0.024 0.014 0.013
Memory MAPE 0.023 0.023 0.027 0.027 0.039 0.022 0.032 0.030 0.019 0.019 0.030 0.019 0.019
Heavy-Workload MSE 0.706 0.673 0.860 0.919 2.420 0.695 1.418 1.182 0.566 0.542 1.114 0.598 0.509
MAE 0.021 0.020 0.023 0.024 0.032 0.019 0.029 0.027 0.018 0.018 0.028 0.018 0.017
MAPE 0.173 0.215 0.181 0.165 0.272 0.190 0.258 0.158 0.139 0.158 0.150 0.153 0.142
Overall MSE 4.647 4.996 4.753 4.809 12.638 6.097 9.954 6.204 4.786 4.812 4.784 6.039 4.758
Alibaba MAE 0.049 0.053 0.051 0.050 0.090 0.060 0.073 0.056 0.050 0.051 0.050 0.054 0.049
CPU MAPE 0.195 0.237 0.194 0.193 0.343 0.247 0.381 0.252 0.178 0.177 0.182 0.176 0.143
Heavy-Workload MSE 16.767 17.548 16.228 17.742 29.493 18.848 32.271 22.226 16.594 16.321 16.405 14.090 11.634
MAE 0.095 0.100 0.096 0.100 0.150 0.105 0.145 0.115 0.097 0.095 0.095 0.087 0.082
(a) Data distribution of different datasets
(b) Time overhead of different methods
(c) MAE of PePNet with different $\gamma$
(d) Overall MAE of PePNet on Alibaba2018-CPU
(e) Heavy-workload MAE of PePNet on Alibaba2018-CPU
(f) Overall MAE of PePNet on Alibaba2018-Memory
(g) Heavy-workload MAE of PePNet on Alibaba2018-Memory
(h) Overall MAE of PePNet on Dinda’s dataset
(i) Heavy-workload MAE of PePNet on Dinda’s dataset
Figure 4: (a) The value of the y-axis is the original workload divided by the maximum workload of its machine. The wider the width of the violin in the figure is, the more workloads are distributed near this value. (b) The figure shows the average consuming time of one forward propagation and backpropagation for one sample. The labels on the histogram are 1000 times of the actual value. Besides, we use the shorthand of Informer, Autoformer and Reformer in the xticks. (c) The figure shows the overall and heavy-workload MAE of PePNet for different $\gamma$ on all the three datasets. In the legend, we use dataset name* to represent the heavy-workload MAE on the specific dataset, while using dataset name to represent the overall MAE on the dataset. (d)-(i) These figures show overall and heavy-workload MAE of PePNet on different datasets with different combination of the length of $X_{in}$ and $p$.
Datasets. We perform experiments on three public datasets: Alibaba’s cluster trace v2018444https://github.com/alibaba/clusterdata, Dinda’s dataset [12] and Server Machine Dataset (SMD) [26]. These datasets are collected by different organizations with different system configurations. Their use can help to verify the generalization of PePNet. We plot the data distribution of different datasets in Fig. 4(a). It is worth noting that different datasets represent different data distributions: long-tailed distribution (Dinda’s dataset), centralized distribution (memory usage of Alibaba2018 and SMD), and relatively even data distribution (CPU usage of Alibaba2018).
  • Dinda’s dataset (long-tailed distribution): Dinda’s dataset is collected from two groups of machines. The first is the Alpha cluster at the Pittsburgh Supercomputing Center (PSC). The second one includes computing servers (Mojave, Sahara), a testbed (manchester1-8), and desktop workstations at Carnegie Mellon University. The workload in Dinda’s dataset shows significant periodicity and sometimes there are large peaks far above average (long-tailed distribution).
  • Alibaba2018-Memory (centralized distribution): Alibaba’s cluster trace is sampled from one of Alibaba’s production clusters, which includes about 4000 machines’ workload in 8 days. Memory usage of some machines in this dataset shows relatively significant periodicity and others are aperiodic. Memory usage is generally stable and is distributed around a certain value (centralized distribution).
  • Alibaba2018-CPU (relatively even distribution): The CPU usage in Alibaba2018 shows more significant periodicity on the heavy workload than the light load. Besides, it is highly variable and is distributed relatively even.
  • SMD dataset (centralized distribution with rare extreme values): SMD is a new 5-week-long dataset, which is collected from a top 10 Internet company. SMD is made up of data from 28 different machines. It consists of 38 features, including CPU utilization, memory utilization, disk I/O, etc. The workload in this dataset shows high periodicity. Besides, most of the workload of this dataset is concentrated around 10-20, while others are extremely big.
We use the first 1,000,000 lines in Alibaba’s dataset and axp0 in Dinda’s dataset. The values in Alibaba2018 are utilization percentage and we uniformly scale it to range [0, 1].
Evaluation metrics. We use three metrics to evaluate PePNet’s performance: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). These metrics are some of the most recognized and widely used ones in time series prediction. There are many influential works using these metrics, such as [11], [5], [23]. Besides, these metrics not only measure the relative error (MAPE) but also the absolute error (MSE, MAE).
III-B Prediction Accuracy
We summarize the performance of all the methods in Tab. IV. We use the above metrics to evaluate the accuracy of heavy-workload forecasting accuracy as well as overall forecasting accuracy. We highlight the highest accuracy in boldface. Besides, if PePNet achieves the highest accuracy, we highlight the best one in baseline with underlines.
As shown in Tab. IV, PePNet achieves the best heavy-workload prediction accuracy as well as overall prediction accuracy on all three workload datasets, with few exceptions. Comparing the prediction accuracy on all three datasets, PePNet works the best on long-tail distributed Dinda’s dataset, while less advantageous on the relatively even-distributed Alibaba2018-CPU dataset.
Long-tail data distribution. PePNet works best in this case. Compared with the best performance among the baselines, PePNet improves MAPE, MSE, and MAE by 17.9%, 24.0%, and 29.5% respectively for overall prediction. For heavy-workload prediction, PePNet improves MAPE, MSE, and MAE by 40.8%, 10.4%, and 38.7% respectively over the best method in the baseline.
Centralized data distribution with rare extreme value. PePNet also has obvious advantages in this case, because the SMD dataset has extremely heavy workload. Compared with the best performance in the baseline, PePNet improves MAPE, MSE and MAE by 14.0%, 25.2% and 26.7% respectively for overall prediction. For heavy-workload prediction, PePNet improves MAPE and MAE by 25.3% and 25.0% respectively.
Centralized data distribution. PePNet is also effective in this case. Compared with the best performance in the baseline, PePNet improves MAPE, MSE and MAE by 33.8%, 13.9% and 9.5% for overall prediction. For heavy-workload prediction, PePNet improves MAPE, MSE and MAE by 15.2%, 24.4% and 8.7% respectively, compared with the best method in the baseline.
Relatively even data distribution. In this case, PePNet maintains high overall prediction accuracy and achieves the best heavy workload prediction performance among neural network-based methods. PePNet improves MAPE, MSE and MAE by 25.7%, 28.3% and 13.9% for heavy-workload prediction over the best method in the baseline.
III-C Time Overhead
We use an Intel(R) Xeon(R) CPU E5-2620 @ 2.10GHz CPU and a K80 GPU to record the time spent on training and inference for all methods. We show the time overhead of all the methods in Fig. 4(b). PePNet brings just an acceptable extra time overhead compared to the most efficient network (GRU) and greatly improves the overall and heavy-workload prediction accuracy, especially on Dinda’s dataset (40.8% for MAPE of heavy-workload prediction).
III-D Hyperparameter Sensitivity
The impact of input length and prediction length. We use grid search to explore the impact of $X^{enc}_{short}$’s length and the impact of $y$’s length on PePNet’s performance. We show the result in Fig. 4(d)-Fig. 4(i). We perform experiments with Cartesian combinations of input lengths from 30 to 80 and prediction lengths from 2 to 12. Generally, the error grows higher gently when predicting length increases. The best input length for different datasets is different. It is found the best input length for Alibaba2018-Memory and Alibaba2018-CPU is 50, while the best input length for Dinda’s dataset is 30.
Dinda’s dataset. For overall workload prediction, MAE rises slowly as prediction length increases and MAE is stable for any input length. For heavy-workload workload prediction, MAE is more stable for prediction length when the input length is longer. As the length of prediction grows, the MAE for overall accuracy increases no more than 0.051 at each input length, while the MAE of heavy workload increases no more than 0.062.
Alibaba2018-Memory. The long-term dependency of Alibaba’s memory usage is weak, thus when the input length is too big, increasing the input length increases the noise level, which could corrupt the performance of PePNet. Thus, when the input length is greater than 50, the MAE of PePNet is unstable. But when the input length is less than 50, the performance of PePNet is stable, and the prediction error only increases slowly with the prediction length. For overall memory usage prediction, the MAE with a prediction length of 12 is only 0.006 larger than the MAE with a prediction length of 2. For heavy-workload memory usage prediction, the MAE with a prediction length of 12 is only 0.012 larger than that with a prediction length of 2.
Alibaba2018-CPU. Long-term dependency is more pronounced in CPU usage. Thus, when we increase the input length, the performance of PePNet is always stable. For overall CPU usage prediction, the MAE with a prediction length of 12 is only 0.016 larger than that with a prediction length of 2. For heavy-workload CPU usage prediction, the MAE with a prediction length of 12 is only 0.034 larger than that with a prediction length of 2.
The impact of $\gamma$. We also explore the impact of $\gamma$ on three datasets, as shown in Fig. 4(c). There are slight fluctuations of overall accuracy for different $\gamma$. Besides, it is consistent with the theoretical analysis in Section.II-C that when the $\gamma$ is smaller the heavy-workload accuracy is higher. It is also intuitive that the heavy-workload MAE of Dinda’s dataset dips most sharply, as Dinda’s dataset is long-tail distributed. The extremely heavy workload in the long tail greatly reduces the prediction accuracy. But as $\gamma$ gets smaller, more attention is put on heavy workload, during the process of training. According to our experimental results, setting $\gamma$ as the value indefinitely approaching zero can lead to the best accuracy. As analyzed in Section.II-C, when $\gamma$ indefinitely approaches zero, Achilles’ loss function only optimizes the part with the highest prediction error. In this situation, PePNet also achieves high prediction accuracy without overall accuracy degradation, which verifies that the Achilles’ loss function helps to avoid overly emphasizing sporadic patterns.
III-E Ablation Experiment
We validate the effect of the Periodicity-Perceived Mechanism and heavy-workload-focused loss function by comparing the performance of PePNet with that of PePNet- and PePNet. We show the performance of these models in Tab. IV and summarize the improvement ratio of PePNet compared with PePNet- and PePNet in Tab. V and Tab. VI respectively. Overall, the Periodicity-Perceived Mechanism contributes more to the prediction accuracy than the Achilles’ Heel loss function. But Achilles’ Heel loss function can improve the accuracy of periodic and aperiodic time series, while the Periodicity-Perceived Mechanism can only promote the accuracy of periodic and lax-periodic data. Furthermore, we also test the prediction accuracy of PePNet on aperiodic data.
The performance of Periodicity-Perceived Mechanism.
Dinda’s dataset and SMD dataset exhibit stricter periodicity than Alibaba2018 does [10]. Thus, the Periodicity-Perceived Mechanism promotes the prediction accuracy most on Dinda’s dataset and SMD dataset. For datasets with lax periodicities, such as memory usage of Alibaba2018, the Periodicity-Perceived Mechanism also promotes overall and heavy-workload prediction accuracy. As for CPU usage of Alibaba2018, which has no significant periodicity on light load but has more significant periodicity on heavy workload, the Periodicity-Perceived Mechanism can promote the heavy-workload prediction accuracy, while the performance of PePNet is about as same as PePNet- on light load. This observation confirms that the Periodicity-Perceived Mechanism has little negative impact on aperiodic data.
TABLE V: Improvement (%) of PePNet over PePNet-.
Overall Heavy-workload
MAPE MSE MAE MAPE MSE MAE
Dinda’s dataset 14.8 36.0 37.1 49.9 -8.7 46.6
SMD dataset 25.7 44.1 30.8 18.7 38.5 21.6
Alibaba2018-Memory 67.6 10.0 9.07 2.53 10.1 2.69
Alibaba2018-CPU -2.1 0.60 0.30 19.4 29.9 15.1
Max. 67.6 44.1 37.1 49.9 38.5 46.6
Min. -2.1 0.59 0.29 2.53 -8.7 2.69
Avg. 26.6 22.7 19.3 22.6 17.5 21.5
The performance of Achilles’ Heel loss function. Our loss function can improve both the overall prediction accuracy and heavy-workload prediction accuracy on all three datasets, compared with PePNet. Except MAPE for heavy-workload prediction accuracy on memory usage of Alibaba2018, all of the evaluation metrics of PePNet are better than that of PePNet. As the Achilles’ Heel Loss function apparently reduces the error for extremely heavy workload in the long tail, which is also proven in Fig. 4(c), the Achilles’ Heel Loss Function improves the accuracy of Dinda’s dataset and SMD dataset more that of Alibaba2018.
TABLE VI: Improvement (%) of PePNet over PePNet.
Overall Heavy-workload
MAPE MSE MAE MAPE MSE MAE
Dinda’s dataset 12.2 39.8 40.4 59.4 16.4 56.6
SMD dataset 19.1 40.9 26.3 12.6 31.0 13.2
Alibaba2018-Memory 39.9 5.28 4.96 -1.3 6.09 0.66
Alibaba2018-CPU 10.3 1.10 2.10 18.9 28.7 13.6
Max. 39.9 40.9 40.4 59.4 31.0 56.6
Min. 10.3 1.12 2.13 -1.3 6.09 0.66
Avg. 20.4 21.8 18.5 22.4 20.6 21.0
TABLE VII: Compare the performance of PePNet on aperiodic and periodic data. The MSE are multiplied by 1000. The MAE are multiplied by 100.
Periodicity Overall Metric Heavy-workload
MAPE MSE MAE MAPE MSE MAE
Ali2018-Memory Periodic 0.024 0.08 0.47 0.018 0.17 0.95
Aperiodic 0.015 0.35 1.35 0.019 0.53 1.78
Both 0.016 0.32 1.27 0.019 0.51 1.74
Ali2018-CPU Periodic 0.133 2.81 3.79 0.117 5.44 5.16
Aperiodic 0.166 8.01 6.87 0.163 16.1 10.2
Both 0.142 4.76 4.94 0.143 12.4 8.21
The performance of PePNet on aperiodic data. There is a major concern about whether PePNet still works well on aperiodic data. To test the prediction accuracy on aperiodic data, we divide the Alibaba2018-CPU and Alibaba2018-Memory by periodicity and collect the prediction accuracy for periodic data and aperiodic data respectively. In Tab. VII, there is prediction accuracy for periodic data and aperiodic data. For convenience, the overall accuracy of both periodic data and aperiodic data is also listed at the bottom. On the whole, the accuracy of periodic data is slightly higher than the overall accuracy, while the accuracy of aperiodic data is slightly lower than the overall. However, there is a strange phenomenon the MAPE of periodic data in Alibaba2018-Memory is bigger than the overall accuracy, while its MAE and MSE are much lower than the overall accuracy. That is because, in the Alibaba2018-Memory dataset, the periodic workload is much smaller than the aperiodic one. Thus, even slight errors in periodic data prediction tend to become much larger after the division in the computation of MAPE.
IV Related Work
We first review existing methods for workload and time series prediction. Since PePNet uses periodic information, we also summarize related work for this topic.
IV-A Workload and Time Series Prediction
Time series prediction aims to predict future time series based on the observed workload data, while workload prediction aims to predict future workload based on observed ones. Present research predicts many different workloads, such as the GPU workload [27], CPU utilization [28], virtual desktop infrastructure pool workload [29], disk states [30], database query arrival rates [31] and so on. Researchers make workload predictions for different aims, such as cloud resource allocation [32], autoscaling [28], etc. Additionally, some works proposed an integrated framework of prediction models and other functional modules. For example, Uncertainty-Aware Heuristic Search (UAHS) [33] combines workload prediction models and Bayesian optimization to model the uncertainty and guide virtual machines (VMs) provision strategy. SimpleTS [34] deploys a framework that automatically selects a proper prediction model for a specific time series. However, these methods only focus on improving overall accuracy while neglecting high load accuracy.
As workload prediction is a branch of time series prediction, it is necessary to review the development of time series prediction. In this domain, several frequently discussed issues garner significant attention. One stream is to improve the prediction accuracy of dynamic real-world time series, especially the accuracy of long input sequences. The solutions for this problem can be roughly divided into four categories: the statistic-based methods [7] [35] [36], the recurrent neural networks [37] [38] [39], the transformer-based methods [23] [24] [11] and the integrated methods [10] [40] [5]. The statistic-based methods are efficient but not robust for highly dynamic real-world time series. The recurrent neural networks improve this aspect but suffer from catastrophic forgetting effects. To conquer this drawback, the integrated methods combine attention mechanism [10] and auto-encoder [14] into the recurrent neural network.  Transformer-based methods make further progress by utilizing transformers. Another time series prediction issue that receives widespread attention is the spatiotemporal forecasting problem. In this field, researchers improve the prediction accuracy by leveraging the spatial and temporal information with GNN, CNN, plugin modules and other neural networks [15] [41] [42] [43]. But these methods manually define the spatio-temporal blocks which capture the temporal and spatial correlations. To overcome this limitation, AutoCTS [44] automatically selects competitive heterogeneous spatiotemporal blocks. But these methods also ignore the under-fitting problem of extreme values. Another research direction is improving the robustness of deep networks dealing with time series. The robustness includes five implications: the robustness for missing data, the robustness for the malicious or limited labels, the generalization of the trained model, the robustness toward noise in time series, and the robustness for the irregular data sampling period. For example, Belkhouja et al. propose a Robust Training for Time-Series (ROTS) to improve deep neural network robustness for generalization and noise [45]; Tan et al. propose Dual-Attention Time-Aware Gated Recurrent Unit (DATA-GRU) to manipulate time series sampled in an irregular period [46]; Zhang et al. propose Multivariate Time Series Classification with Attentional Prototypical Network (TapNet) to deal with time series with limited labels [47]; Tang et al. propose network modeling local and global temporal dynamics (LGnet) to deal with a data-missing problem in time series [48]; Luo et al. propose $E^{2}GAN$, which uses adversarial learning to deal with data missing problem [49]; Yao et al. propose Spatial-Temporal Dynamic Network (STDN) to deal with lax-strict periodicity caused by noise [16]; Luo et al. propose Uncertainty-Aware Heuristic Search (UAHS) to deal with uncertain prediction error caused by noise [33]. However, none of these methods focus on dealing with the prediction accuracy of heavy workload.
IV-B Periodicity Information Extraction
In the field of traffic forecasting, there is also significant periodicity and we summarize some recent popular methods below. There are several mechanisms to fuse periodic information. For example, Guo et al. propose a novel attention-based spatial-temporal graph convolution network [13]; Lv et al. use a fully connected layer and take advantage of both CNN and RNN to deal with periodic time series [15]; Chen et al. propose Hop Res-RGNN to deal with periodic patterns [14]. However, these methods do not consider the lax periodicity of time series. Yao et al. propose an attention mechanism to tackle periodic shift [16]. These studies use the priori knowledge of the daily and weekly periodicity of traffic loads. However, in the scenario of service load prediction, there is usually no priori knowledge about the periodicity. Besides, the periodicity of the time series in these studies is fixed, while the periodicities of workload for different machines are variable.
V Conclusion
In this paper, we study the problem of improving overall workload prediction accuracy as well as heavy-workload prediction accuracy and propose PePNet. PePNet makes use of short-term dependent information, long-term tendency information and periodic information. Within PePNet, we propose two mechanisms for better prediction accuracy of workloads: (1) a Periodicity-Perceived Mechanism to guide heavy-workload prediction, which can mine the periodic information and adaptively fuse periodic information for periodic, lax periodic and aperiodic time series; and (2) an Achilles’ Heel Loss Function to offset the negative effect of data imbalance. We also provide theoretical support for the above design by: (1) providing a theoretically proven error bound of periodic information extracted by Periodicity-Perceived Mechanism; and (2) providing an automatic hyperparameter determination method for periodicity threshold $\mathcal{T}$. Compared with existing methods, extensive experiments conducted on Alibaba2018, Dinda’s dataset and SMD dataset demonstrate that PePNet improves MSE for overall workload prediction by 11.8% on average. Especially, PePNet improves MSE for heavy workload prediction by 21.0% on average.
Acknowledgments
This work was supported by the National Science Foundation of China under Grants 62125206.